The sum of the ages of a man and his wife is 73 years. Eight years ago the husband was twice as old as his wife. How old are they now?
Answer by ankor@dixie-net.com(22740)

The sum of the ages of a man and his wife is 73 years.
m + w = 73
:
Eight years ago the husband was twice as old as his wife.
m - 8 = 2(w-8)
m - 8 = 2w - 16
m - 2w = - 16 + 8
m - 2w = - 8
subtract from the 1st equation
m + w = 73
m - 2w = -8
-----------------subtraction eliminates m, find w
3w = 81
w = 81/3
w = 27 yrs is the wife's age
:
I'll let you find the man's age, check solutions in both equations


How old are they now?